40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 21 18:08:35 2021, yehonathan, Update, Loss Measurement, Loss measurement 
    Reply  Thu Jul 22 16:06:10 2021, Paco, Update, Loss Measurement, Loss measurement 
       Reply  Sun Jul 25 20:41:47 2021, rana, Update, Loss Measurement, Loss measurement 
          Reply  Mon Jul 26 17:34:23 2021, Paco, Update, Loss Measurement, Loss measurement LossMeasurement_RawData.pdfYARM_loss_stats.pdfXARM_loss_stats.pdf
Message ID: 16253     Entry time: Wed Jul 21 18:08:35 2021     Reply to this: 16254
Author: yehonathan 
Type: Update 
Category: Loss Measurement 
Subject: Loss measurement 

{Gautam, Yehonathan, Anchal, Paco}

We prepared for the loss measurement using DC reflection method. We did the following changes:

1. REFL55_Q was disconnected and replaced with MC_T cable coming from the PD on the MC2 table. The cable has a red tag on it. Consequently we lost the AS beam. We realigned the optics and regained arm locks. The spot on the AS QPD had to be corrected.

2. We tried using AS55 as the PD for the DC measurement but we got ratios of ~ 0.97 which implies losses of more than 100 ppm. We decided to go with the traditional PD520 used for these measurements in the past.

3. We placed the PD520 used for loss measurements in front of the AS55 PD and optimized its position.

4. AS110 cable was disconnected from the PD and connected to PD520 to be used as the loss measurement cable.

5. In 1Y2 rack, AS110 PD cable was disconnected, REFL55_I was disconnected and AS110 cable was connected to REFL55_I channel.

So for the test, the MC transmission was measured at REFL55_Q and the AS DC was measured at REFL55_I.

We used the scripts/lossmap_scripts/armLoss/measArmLoss.py script. Note that this script assumes that you begin with the arm locked.

We are leaving the IFO in the configuration described above overnight and we plan to measure the XARM loss early AM. After which we shall restore the affected electrical and optical paths.


We ran the /scripts/lossmap_scripts/armLoss/measureArmLoss.py script in pianosa with 25 repetitions and a 30 s "duty cycle" (wait time) for the Y arm. Preliminary results give an estimated individual arm loss of ~ 30 ppm (on both X/Y arms) but we will provide a better estimate with this measurement. 

ELOG V3.1.3-